OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x^3 + 2x^2 – 24x ?

OpenStudy (hoblos):

take x as a common factor then solve the quadratic

OpenStudy (diyadiya):

x^3 + 2x^2 – 24x=0 x(x^2+2x-24)=0 x(x+6)(x-4)=0 x=0 x=-6 x=4

OpenStudy (turingtest):

take note of how diyadiya solved the problem, that is a useful way to approach it.